### v2

# Gauge fields induced by curved spacetime

Abstract
I found an extended duality (triality) between Dirac fermions in periodic spacetime metrics, nonrelativistic fermions in gauge fields (e.g., Harper-Hofstadter model), and in periodic scalar fields on a lattice (e.g., Aubry-André model). This indicates an unexpected equivalence between spacetime metrics, gauge fields, and scalar fields on the lattice, which is understood as different physical representations of the same mathematical object, the quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$. This quantum group is generated by the exponentiation of two canonical conjugate operators, namely a linear combination of position and momentum (periodic spacetime metrics), the two components of the gauge invariant momentum (gauge fields), position and momentum (periodic scalar fields). Hence, on a lattice, Dirac fermions in a periodic spacetime metric are equivalent to nonrelativistic fermions in a periodic scalar field after a proper canonical transformation. The three lattice Hamiltonians (periodic spacetime metric, Harper-Hofstadter, and Aubry-André) share the same properties, namely fractal phase diagrams, self-similarity, topological invariants, flat bands, and topological quantized current in the incommensurate regimes.
Gravity remains the lone outlier among the fundamental forces, defying all efforts to fit it into a coherent quantum theory — a challenge that continues to puzzle physicists to this day. However, in quantum field theory on curved spacetime[1], one usually ignores these issues and describes, e.g., fermions acting in the presence of gauge fields and within the backdrop of a spacetime metric. Yet, even in this approach, gravity seems to have a special role, being the deus ex machina that bends or reshapes the stage where gauge fields and fermions perform. Hence, from the viewpoint of a fermion, a gauge field, a scalar field, and a spacetime metric are very different objects.
In the light of the above, the result obtained here is surprising: On a discrete lattice[2], a periodic spacetime metric is equivalent to a gauge field and to a periodic scalar field. More precisely, I found that the Hamiltonian of a Dirac fermion in a curved spacetime[3, 4] described by a periodic spacetime metric is equivalent to the Hamiltonian of a nonrelativistic fermion in a gauge field (Harper-Hofstadter model[5, 6]) or in a periodic scalar field (Aubry-André model[7]) when regularized on the lattice. While the duality between nonrelativistic lattice fermions in gauge fields (Harper-Hofstadter) and periodic scalar fields (Aubry-André model) is well known, the connection with Dirac fermions in curved periodic spacetime on a lattice is unexpected: In essence, these three different models are all different physical representations of the same quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$, generated by the exponentiation of two canonical conjugate operators. These operators are the two components of the gauge invariant momentum (Harper-Hofstadter model), position and momentum (Aubry-André model), and a linear combination of position and momentum (periodic spacetime metrics). Hence, on a lattice, Dirac fermions in a periodic spacetime metric are obtained via a canonical transformation of nonrelativistic fermions in a periodic scalar field. This triality (i.e., nontrivial equivalence between three different models) mandates that all three models share the same properties, e.g., fractal phase diagrams, self-similarity, topological invariants, flat bands and topological quantized current in the incommensurate regimes, as I will shortly illustrate. Hamiltonians describing Dirac fermions in a curved spacetime regularized on a lattice are often considered in condensed matter[8, 9, 10, 11, 12, 13, 14]. These results may open new venues for studying analog gravity.
Nonrelativistic fermions on a 2D discrete lattice in a gauge field are described by the so-called Harper-Hofstadter Hamiltonian. To derive this Hamiltonian, one can start from charged fermions in a gauge field and then treat the discrete 2D lattice (described by a periodic potential) as a perturbation on the Landau levels (as in Ref. 15). Conversely, one can start with a so-called "tight-binding" model describing charged fermions in a periodic potential and treating the gauge field as a perturbation (as in Refs. 5, 6). These two approaches lead to the same result, namely
[…]
where the two terms are the contributions of the gauge-invariant momentum $\hat{\boldsymbol{\pi}}=\hat{\mathbf{p}}-\hat{\mathbf{A}}$ regularized on the lattice, where $\hat{\mathbf{A}}=(-\omega_{y}\hat{y},\omega_{x}\hat{x})$ of a uniform field in Coulomb gauge. The two components of the gauge-invariant momentum are canonical conjugate operators $[\hat{\pi}_{x},\hat{\pi}_{y}]=\mathrm{i}\omega$ with $\omega=\omega_{x}+\omega_{y}$ equal to the flux of the field per unit cell.
On the other hand, nonrelativistic fermions on a 1D discrete lattice in a periodic scalar field are described by the so-called Aubry-André model[7], which can be written as
$\mathcal{H}_{\text{AA}}=2\cos{\left(\omega\hat{x}+\phi\right)}+2\cos{\hat{p}},$ (2)
where the first term is the contribution of the scalar field $\propto\cos{\left(\omega{x}+\phi\right)}$ ), and the second term is the particle momentum regularized on the lattice.
The duality between $\mathcal{H}_{\text{HH}}$ and $\mathcal{H}_{\text{AA}}$ is evident: both Hamiltonians can be written as
[…]
where $\hat{X},\hat{Y}$ are the canonical conjugate operators $[\hat{X},\hat{Y}]=\mathrm{i}\omega$. Thus, taking $\hat{X}=\hat{p}_{x}$ and $\hat{Y}=\hat{p}_{y}$ yields the Harper-Hofstadter model in Eq. 1; taking $\hat{X}=\omega\hat{x}+\phi$ and $\hat{Y}=\hat{p}$ yields the Aubry-André model in Eq. 2. The quantity $\omega=-\mathrm{i}[\hat{X},\hat{Y}]$ describes the flux of the gauge field per unit cell in the Harper-Hofstadter model and the wavenumber of the periodic scalar field in the Aubry-André model. The phase shift $\phi$ of the periodic scalar field in the Aubry-André model plays the same role as one of the components of the momentum in the Harper-Hofstadter model and, for this reason, can be regarded as a "synthetic" dimension[16]. In the quantum geometry language[17, 18, 19, 20, 21, 22, 23], the Hamiltonian Eq. 3 is $\mathcal{H}_{XY}=\hat{T}_{X}+\hat{T}_{X}^{\dagger}+\hat{T}_{Y}+\hat{T}_{Y}^{\dagger}$ where $\hat{T}_{X}=\mathrm{e}^{\mathrm{i}\hat{X}}$ and $\hat{T}_{Y}=\mathrm{e}^{\mathrm{i}\hat{Y}}$ are the so-called magnetic translation operators, satisfying the algebraic equations $\hat{T}_{X}\hat{T}_{Y}=\mathrm{e}^{\mathrm{i}\omega}\hat{T}_{Y}\hat{T}_{X}$, which are a quantum "deformation" of the usual algebra of spatial translations, which is commutative $\hat{T}_{X}\hat{T}_{Y}=\hat{T}_{Y}\hat{T}_{X}$. These operators $\hat{T}_{X}$ and $\hat{T}_{Y}$ coincide with two of the generators of the quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ obtained as a $q$-deformation of the enveloping algebra $U(\mathfrak{sl}_{2})$ of the Lie algebra $\mathfrak{sl}_{2}$, with $q^{2}=\mathrm{e}^{\mathrm{i}\omega}$.
However, the story does not necessarily end here: In principle, one can consider any set of conjugate variables $\hat{X},\hat{Y}$ in Eq. 3. An interesting choice is $\hat{X}=\frac{1}{2}(\omega\hat{x}+\phi)-\hat{p}$ and $\hat{Y}=\frac{1}{2}(\omega\hat{x}+\phi)+\hat{p}$ giving again $[\hat{X},\hat{Y}]=\mathrm{i}\omega$ and
$\mathcal{H}_{\text{CS}}=4\cos{\left(\frac{1}{2}(\omega\hat{x}+\phi)\right)}\cos{\hat{p}},$ (4)
where the subscript stands for curved spacetime, as clarified hereafter. Regularizing on the lattice
$\mathcal{H}_{\text{CS}}=\sum\limits_{n}2\cos\left(\frac{1}{2}(\omega n+\phi)\right)\Big{(}\ket{n+1}\bra{n}+\ket{n-1}\bra{n}\Big{)}.$ (5)
Note that this Hamiltonian coincides with the Su-Schrieffer-Bardeen model for polyacetylene chains[24] for $\omega=2\pi$, and hence is a generalization of that model to arbitrary spatial wavelength $\omega\neq 2\pi$.
Hence, not only the Hamiltonians $\mathcal{H}_{\text{HH}}$ and $\mathcal{H}_{\text{AA}}$ are dual, but they are also dual to the Hamiltonian $\mathcal{H}_{\text{CS}}$ in Eq. 4, since all three Hamiltonians are formally expressed as in Eq. 3. In particular the Hamiltonian $\mathcal{H}_{\text{AA}}$ is related to the Hamiltonian in Eq. 4 by a canonical transformation $\hat{X}\to\frac{1}{2}\hat{X}-\hat{Y}$, $\hat{Y}\to\frac{1}{2}\hat{X}+\hat{Y}$. Hence, the models $\mathcal{H}_{\text{HH}}$, $\mathcal{H}_{\text{AA}}$, and $\mathcal{H}_{\text{CS}}$ are dual one to the other: I refer to this property as a triality.
To gain an intuition about the physical meaning of the Hamiltonian $\mathcal{H}_{\text{CS}}$, notice that in position basis, Eq. 5 describes a fermion hopping back and forth with spatially modulated hopping amplitudes. This spatial modulation intuitively suggests the presence of a warped or deformed spacetime metric. This is because the covariant derivatives of a quantum field equation in curved spacetime typically correspond to space-dependent hopping amplitudes when regularized on the lattice. Indeed, this intuition is correct, as I will show hereafter.
Consider a massless Dirac fermion in 1+1D curved spacetime[3, 4]
$\left[\mathrm{i}\gamma^{a}e_{a}{}^{\mu}\partial_{\mu}+\frac{\mathrm{i}}{2}\gamma^{a}\frac{1}{\sqrt{-g}}\partial_{\mu}(\sqrt{-g}\,e_{a}{}^{\mu})\right]\psi=0,$ (6)
where $\psi$ is a spinor, $\gamma^{\mu}$ the flat spacetime Dirac gamma matrices satisfying $\{\gamma^{\mu},\gamma^{\nu}\}=2\eta^{\mu\nu}$ with $\eta_{\mu\nu}=\operatorname{diag}{(1,-1)}$ the Minkowski metric, $\sqrt{-g}$ the square root of the determinant of the metric, and the zweibein $g_{\mu\nu}=e^{a}{}_{\mu}e^{b}{}_{\nu}\eta_{ab}$, and $\eta_{ab}=e_{a}{}^{\mu}e_{b}{}^{\nu}g_{\mu\nu}$, with $g^{\mu\nu}=(g_{\mu\nu})^{-1}$. In the Weyl representation $\gamma^{0}=\sigma_{x}$ and $\gamma^{1}=\mathrm{i}\sigma_{y}$. Consider the metric
[…]
which yields $g_{00}=\alpha(x)^{2}$ and $g_{11}=-1$, $e_{0}{}^{0}=\alpha(x)^{-1}$ and $e_{1}{}^{1}=1$, and $\sqrt{-g}=\alpha(x)$. Separating time and space components
$\mathrm{i}\partial_{0}\psi=H\psi=-{\mathrm{i}\sqrt{\alpha(x)}}\gamma_{0}\gamma^{1}\partial_{1}\left(\sqrt{\alpha(x)}\psi\right),$ (8)
which describes the time evolution of the spinor field, with $H$ the Hamiltonian density. The corresponding Hamiltonian can be regularized on a discrete lattice using
$\partial_{1}\left(\sqrt{\alpha(x)}\psi\right)\approx\frac{1}{2}\left(\sqrt{\alpha_{n+1}}\psi_{n+1}-\sqrt{\alpha_{n-1}}\psi_{n-1}\right),$ (9)
where $\alpha_{n}=\alpha(n)$ is the discretized metric, which yields
$\mathcal{H}=\mathrm{i}\sum\limits_{n}t_{n}\psi_{n+1}^{\dagger}\gamma_{0}\gamma^{1}\psi_{n}-t_{n}\psi_{n}^{\dagger}\gamma_{0}\gamma^{1}\psi_{n+1},$ (10)
where $t_{n}=\frac{1}{2}\sqrt{\alpha_{n}\alpha_{n+1}}$. This Hamiltonian is hermitian. The gauge transformation $\psi_{n}\to\mathrm{i}^{n}\psi_{n}$ yields
$\mathcal{H}=\sum\limits_{n}t_{n}\left(\psi_{n+1}^{\dagger}\gamma_{0}\gamma^{1}\psi_{n}+\psi_{n}^{\dagger}\gamma_{0}\gamma^{1}\psi_{n+1}\right).$ (11)
Now, consider the metric
${\alpha_{n}}=\frac{1}{C^{(-1)^{n}}}\prod_{m=1}^{n-1}\left[\cos^{2}\left(\frac{1}{2}(\omega m+\phi)\right)\right]^{(-1)^{m+n+1}},$ (12)
for $n\geq 1$ and with $C$ an arbitrary number defining the boundary value $\alpha_{1}=C$. In this metric $2t_{n}=\sqrt{{\alpha_{n}\alpha_{n+1}}}=|\cos{\left(\frac{1}{2}(\omega n+\phi)\right)}|$, and therefore the Hamiltonian $\mathcal{H}$ in Eq. 11 becomes equivalent to the Hamiltonian $\mathcal{H}_{\text{CS}}$ in Eq. 4 describing massless Dirac fermion in a periodic spacetime metric on a 1D lattice.
When the wavelength is commensurate with the lattice, i.e., when $\omega=2\pi p/q$ with $p,q$ coprimes, the hopping amplitudes are periodic. Conversely, when the wavelength is incommensurate, the hopping amplitudes are quasiperiodic. Note that the metric on the lattice $\alpha_{n}$ diverges for values of the phase $\omega m+\phi=\pm\pi$ on lattice sites where $m+n$ is even. The values of the phase where the metric is singular form a set $\{\phi=\pm\pi-\omega m\mod 2\pi,m\in\mathbb{Z}\}\subset[0,2\pi]$. In the commensurate case, the cardinality of the set is $q$ or $2q$ depending on whether $q$ is even or odd, i.e., there is a finite number of values ($q$ or $2q$) of the phase $\phi$ where the metric is singular. In the incommensurate case, the cardinality of the set is $\aleph_{0}$, i.e., there is an infinite (and countable) number of values of the phase where the metric is singular. Note that, even in cases where the metric is singular, the hopping amplitudes $t_{n}\propto\sqrt{\alpha_{n}\alpha_{n+1}}$ remain finite and well-defined. Note also that $\alpha_{n}$ is not, in general, periodic, in contrast with the hopping amplitudes $t_{n}$ that are always periodic in the commensurate regime or quasiperiodic in the incommensurate regime. In particular, one can show that $\alpha_{n}$ is periodic on the lattice with period $2q$ in the commensurate regime only for $q$ odd [^1] See Supplemental Material. However, for the sake of simplicity, I refer to the metric as being periodic or quasiperiodic as long as the hopping amplitudes on the lattice are periodic or quasiperiodic.
The lattice Hamiltonians $\mathcal{H}_{\text{HH}}$ and $\mathcal{H}_{\text{AA}}$ describing a massive nonrelativistic fermion respectively in a gauge field and in a periodic scalar field, together with the Hamiltonian $\mathcal{H}_{\text{CS}}$ describing a massless Dirac fermion in a periodic spacetime metric are dual one to the other, being three different incarnations of the general Hamiltonian $\mathcal{H}_{XY}$ in Eq. 3. This triality is summarized in Fig. 1. Note that the role of the quantity $\omega$ is played by the flux of the gauge field per unit cell, the wavelength of the periodic scalar field, and the wavelength of the spacetime metric, respectively, for $\mathcal{H}_{\text{HH}}$, $\mathcal{H}_{\text{AA}}$, and $\mathcal{H}_{\text{CS}}$.
Figure 1: Triality between gauge fields, periodic scalar fields, and curved spacetime metrics on finite lattices. (left) The Hamiltonian $\mathcal{H}_{\text{AA}}$ of a nonrelativistic charged fermion on a 1D lattice in a periodic scalar field with wavelength $\omega$ and phase shift $\phi$ (Aubry–André model) is equivalent under the transformation $(\omega\hat{x},\hat{p})\to(\hat{\pi}_{x},\hat{\pi}_{y})$ to (right) the Hamiltonian $\mathcal{H}_{\text{HH}}$ of a nonrelativistic charged fermion on a 2D square lattice in a gauge field with flux $\omega$ per unit cell (Harper-Hofstadter model). These two Hamiltonians are equivalent under a canonical transformation to (bottom) the Hamiltonian $\mathcal{H}_{\text{CS}}$ of a massless relativistic fermion in a periodic spacetime metric with wavelength $\omega$. These three Hamiltonians $\mathcal{H}_{\text{AA}}$, $\mathcal{H}_{\text{HH}}$, and $\mathcal{H}_{\text{CS}}$ can be all written as $\mathcal{H}_{XY}=2\cos\hat{X}+2\cos\hat{Y}$ where $[\hat{X},\hat{Y}]=\mathrm{i}\omega$ and with the two canonical conjugate varaibles being respectively (left) $\hat{X},\hat{Y}=\omega\hat{x},\hat{p}$, (right) $\hat{X},\hat{Y}=\hat{\pi}_{x},\hat{\pi}_{y}$, and (bottom) $\hat{X},\hat{Y}=\frac{1}{2}\omega\hat{x}+\hat{p},\frac{1}{2}\omega\hat{x}-\hat{p}$.
Figure 2: Energy spectra of the Hamiltonian $\mathcal{H}_{\text{CS}}$ of a massless Dirac fermion in a periodic spacetime metric on a lattice of $N=210$ sites. (a) Energy spectra as a function of $p/q$ with periodic boundary conditions and $\phi=0$. (b) Energy spectra for $p/q=1/3$ as a function of the phase $\phi$ with open boundary conditions. (c) The metric $\alpha_{n}$ for $p/q=1/3$ and $\phi=0$ compared with $\cos{\left(\frac{1}{2}(\omega n+\phi)\right)}$ ) (dashed) as a function of the lattice site. (d) Wavefunctions of the two edge modes in the first band gap (from below) for $p/q=1/3$ and $\phi=0$. (e) Energy spectra for $\omega\approx 2\pi(\Phi-1)$ (with $\Phi$ the golden ratio), as a function of the phase $\phi$ with periodic boundary conditions on a lattice of $N=233$ sites. (f) The metric $\alpha_{n}$ for $\omega\approx 2\pi(\Phi-1)$ and $\phi=0$ as a function of the lattice site. (g) Wavefunctions of two bulk modes for $\omega\approx 2\pi(\Phi-1)$ and $\phi=0$.
The triality mandates that every property valid for any of the three Hamiltonians is valid, mutatis mutandis, for all the other two. I will now go through some of these properties.
If the wavelength is commensurate to the lattice, the spectrum of the Hamiltonian $\mathcal{H}_{\text{CS}}$ in Eq. 4 has exactly $q$ energy bands spanning the phase $\phi$ and momentum $k$. Each of the open gaps $j$ is labeled by a nonzero topological invariant $c\in\mathbb{Z}$ given by the Chern number[27, 15, 28] $|c|<q/2$ satisfying the diophantine equation[29, 30, 31] $pc=j\mod q$. These nontrivial gaps exhibit $|c|$ topologically protected edge modes localized at the boundaries when considering a finite lattice of $N=mq$ sites with open boundary conditions, as it follows from the bulk-boundary correspondence[32].
For incommensurate wavelengths $\omega$, i.e., when $\omega/2\pi$ is an irrational number $\in\mathbb{R}-\mathbb{Q}$, the energy spectrum is a Cantor set[33], and the dispersion of the Bloch bands becomes flat[34, 35, 36]. Hence, the periodic spacetime metric induces a fractal phase diagram spanned by the wavelength $\omega$ with infinitely many nontrivial gaps, as in the Hofstadter butterfly[6].
This phase diagram shows a remarkable self-similarity property[21, 22]: It is invariant under the action of the duality transformation $p/q\to q/p,E\to\widetilde{E}$, with $\widetilde{E}$ given by some unknown function and with $p/q$ defined modulo 1 (or equivalently $\omega=2\pi p/q$ defined modulo $2\pi$). This property has been related to the notion of modular double[37] and, by extension, to the Langlands duality of quantum groups[22]. In the Harper-Hofstadter model, the two branches of the duality transformation $p/q\to q/p$ correspond to two opposite limits, obtained by considering the periodic potential as a perturbation on the Landau levels, or by considering the gauge field as a perturbation on the tight-binding model describing particles trapped in a periodic potential, as already noted in Ref. 15. In the Aubry-André model and in the curved spacetime model in Eq. 4, this duality transformation corresponds to switching the wavelength $\omega\to(2\pi)^{2}/\omega$ of the periodic scalar field and of the periodic spacetime metric, respectively. The spectrum is given by the roots of the characteristic polynomial expressed by the Chambers relation[38]
$\det(\mathcal{H}_{\text{CS}}-E)=\det(\mathcal{H}_{\text{AA}}-E)=$
$f_{p/q}(E)-2(-1)^{q}\left(\cos(qk)+\cos(q\phi)\right),$ (13)
where with abuse of notation, $\mathcal{H}_{\text{CS}}$, $\mathcal{H}_{\text{AA}}$ denote here the matrices expressing the corresponding Hamiltonians in the basis of plane waves $\mathrm{e}^{\mathrm{i}kn}$ on the lattice, with the function $f_{p/q}(E)$ given by a polynomial in $E$ (see Ref. 23). The spectrum is given by the roots of the characteristic polynomial $\mathcal{P}(E)=\det(\mathcal{H}_{\text{CS}}-E)=\det(\mathcal{H}_{\text{AA}}-E)$, where with abuse of notation, $\mathcal{H}_{\text{CS}}$, $\mathcal{H}_{\text{AA}}$ denote here the matrices expressing the Hamiltonians in the basis of plane waves $\mathrm{e}^{\mathrm{i}kn}$. This polynomial is given by the Chambers relation[38] $\mathcal{P}(E)=f_{p/q}(E)-2(-1)^{q}\left(\cos(qk)+\cos(q\phi)\right)$, with the function $f_{p/q}(E)$ given by a polynomial in $E$ (see Ref. 23). An analogous expression for $\det(\mathcal{H}_{\text{HH}}-E)$ is obtained by taking $k\to k_{x}$ and $\phi\to k_{y}$. Remarkably $f_{p/q}(E)=f_{q/p}(\widetilde{E})$ under the duality transformation $p/q\to q/p$[21, 22].
In the Aubry-André model, nonzero Chern numbers correspond to the quantization of the charge transport and to the so-called topological charge pump[27]. The same can be said in the spacetime metric model Hamiltonian $\mathcal{H}_{\text{CS}}$: During an adiabatic evolution of the phase $\phi=0$ to $2\pi$, the center of mass of $j$ particles on the lattice describes a trajectory that corresponds to a finite number of particles (i.e., amount of charge) transferred through the lattice, and this number is an integer equal to the Chern number $c$. In the incommensurate case, moreover, not only the charge transferred is quantized, but also the corresponding current: Precisely, the pumped charge becomes linear in time, and the instantaneous current (i.e., the time derivative of the charge pumped) becomes constant, being topologically quantized and equal to the Chern number[36]. Hence, one obtains a quantization of the current induced by the spacetime metric on a lattice.
Equation 4 is, in essence, a discrete tight-binding Hamiltonian with lattice-dependent hopping amplitudes. As such, there are several physical systems that simulate this model, e.g., arrays of lattices of atoms deposited on a surface[39, 40], arrays of quantum dots[41], cold atoms in optical lattices[42, 43, 44, 45, 46, 47], photonic crystals[16], superconducting quantum circuits[48, 49, 50], topologically nontrivial stripes[51], and exciton-polariton condensates in artificial lattices[52]. Controlling the hopping amplitudes corresponds to controlling the distances and overlap amplitudes between contiguous localized states on the lattice.
Figure 2(a) show the energy spectra of the Hamiltonian $\mathcal{H}_{\text{CS}}$ (Hofstadter butterfly) describing massless Dirac fermions in curved spacetime on a lattice as a function of the wavelength $\omega$ of the periodic spacetime metric. Figures 2(b) to 2(d) show the energy spectrum and edge modes calculated with open boundary conditions for $p/q=1/3$ and the corresponding periodic spacetime metric in Eq. 12. Figures 2(e) to 2(g) show the energy spectrum and some of the bulk modes calculated with periodic boundary conditions and the corresponding quasiperiodic spacetime metric in Eq. 12 for a wavelength incommensurate with the lattice, specifically $\omega=2\pi(\Phi-1)$, where $\Phi$ is the golden ratio. Note that the dispersion of the energy bands in the synthetic dimension $\phi$ is flattened, which corresponds to energy modes delocalized on the entire lattice. Note also that the spacetime metric is not periodic anymore and shows large amplitude oscillations.
In conclusion, I extended the duality between the Harper-Hofstadter model and the Aubry-André model, describing lattice fermions in a gauge field and in a periodic scale field, to a triality between these two models and a model describing lattice Dirac fermions in a periodic spacetime metric. This unveils an unexpected equivalence between spacetime metrics, gauge fields, and scalar fields on the lattice. This triality is the equivalence of these three different models being different physical representations of the same quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$. This quantum group is generated by the exponentiation of two canonical conjugate operators, namely the two components of the gauge invariant momentum (Harper-Hofstadter model), position and momentum (Aubry-André model), and a linear combination of position and momentum (periodic spacetime metric). Hence, on the lattice, a Dirac fermion in a periodic spacetime metric is equivalent to a nonrelativistic fermion in a periodic scalar field after a proper canonical transformation. I thus derived several properties of Dirac fermions on a lattice in a periodic spacetime metric by applying the triality, namely, the fractality of the phase diagram, self-similarity properties, topological invariants, flat bands and topological quantized current in the incommensurate regimes.
Acknowledgements.
This work is supported by the Japan Science and Technology Agency (JST) of the Ministry of Education, Culture, Sports, Science and Technology (MEXT), JST CREST Grant. No. JPMJCR19T2, the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Early-Career Scientists Grants No. JP23K13028 and No. JP20K14375.